Calculate the area inside the curve given byr = 5sin and outside the curve given by r = 2 +sine al 21 e) None of these
Answer #1

Sol Griven curves 8=5 sin e 8=2+ sino. Consider the sketches feasible region r-5 sind intersection points when s sino = 2 +a) Required area ੧ p=6 4 f (6 sino-sino-1) do 3 . 8 - & 2 - 2 t Casa & /6 =4 3 Tv 3, * 0---*ਤੂੰ ਹੀ +) 3 \S -) 0 4 . 3 ਬੱਣ +